# Deep Learning for Subspace Regression

### v1

@@ 2.2 Examples of subspace regression problem
When equation (5) is discretised with finite difference or finite element method, it reduces to linear problem $A\phi=f$, where $A$ is large sparse matrix and $\phi,f$ are discretised solution and right-hand side of (5). To exploit sparsity of $A$ one can solve linear equation with relaxation method. General relaxation method split matrix additively $A=D+C$, where $D$ is regular with known inverse [27]. Given the split, if iteration scheme $x^{n+1}=x^{n}+D^{-1}\left(b-Ax^{n}\right)$ is convergent, steady state is exact solution $x=A^{-1}b$. Convergence is linear and its rate is defined by spectral radius of error propagation matrix $I-D^{-1}A$. To improve convergence rate, one can augment iterative method with coarse-grid correction [33]. This techniques allows one to remove influence of leading subspace $V$ of $I-D^{-1}A$ by solving small reduced linear system for error equation $V^{\top}AVe=r$, where $e$ and $r$ are error and residual in the subspace $\mathcal{S}(V)$. Naturally, subspace regression (1) for this problem approximates the mapping $A\rightarrow\mathcal{S}(V)$ or $k(x)\rightarrow\mathcal{S}(V)$ for linear systems resulting from equation (5).
Deflation for conjugate gradient. Krylov subspace methods provide a more systematic way to solve large sparse linear systems [27]. For linear system with symmetric positive definite matrix $A$ resulting from discretisation of equation (5), the method of choice is conjugate gradient (CG) [15]. Similar to other Krylov methods, on step $r$, CG identify optimal solution within Krylov subspace $\mathcal{K}_{r}=\text{span}\left\{b,Ab,\dots,A^{r-1}b\right\}$, where span refers to the subspace formed by linear combinations of vectors in the set. Since powers of $A$ are involved, the most readily available vectors are from the subspaces with large eigenvalues [28]. To improve convergence of method it is reasonably to include eigenspaces $V$ with small eigenvalues to the approximation space $\mathcal{K}_{r}$. The resulting method is deflated CG and the approximation space is $\mathcal{K}_{r}\cup\mathcal{S}(V)$ [29]. In analogy with previous example, the subspace regression problem considered approximates $A\rightarrow\mathcal{S}(V)$ or $k(x)\rightarrow\mathcal{S}(V)$, but this time $V$ spans eigenspaces with small eigenvalues of matrix $A$.
@@ 3.1 Loss functions
    Loss function $L_{1}(A,B)=p-\left\|Q_{B}^{\top}Q_{A}\right\|_{F}^{2}$ satisfies requirements (2), where $A=Q_{A}R_{A},B=Q_{B}R_{B}$ are QR decompositions of $A$ and $B$, $\left\|\cdot\right\|_{F}$ is Frobenius norm.
@@
    Let $z\in\mathbb{R}^{k}$ be a random variable with zero mean and identity covariance matrix. Loss functions $L_{2}(A,B;z)=\min_{u}\left\|Au-Q_{B}z\right\|_{2}^{2}$ does not depend on the choice of $A$ from $\lceil A\rceil$, where $B=Q_{B}R_{B}$ is QR decomposition.
@@ 4.1 Eigenspace prediction
Loss without $QR$ scales much better for larger subspace sizes. In Figure 1(a) we demonstrate wall clock training time for $L_{1}(A,B)$ and $L_{2}(A,B;z)$ (least squares problem is solved with normal equation) per epoch on the same hardware for the same FFNO architectures. For small subspace sizes the training time the methods are roughly on par, but with the increase of subspace size QR starts to drastically slow down training with $L_{1}$ loss.
@@
Loss $L_{2}(A,B;z)$ may become unstable. From results summarised in Figure 1(a) one can assume that $L_{2}(A,B;z)$ is always preferable. Results for $D=3$ problem (elliptic eigenproblem, grid $30\times 30\times 30$, prediction of first $3$ eigenvectors) summarised in Table 2 reveal a more nuanced picture. Loss $L_{2}(A,B;z)$ clearly performs worse than $L_{1}(A,B)$ and even fails for subspace size $24$. The reason for that is numerical instability of solvers based on the normal equation. To stabilize $L_{2}(A,B;z)$ we apply Cholesky-QR2 [37]. The results for stabilised loss shows that accuracy becomes comparable to $L_{1}(A,B)$ and even slightly better for larger subspace dimensions.
@@ 4.3 Iterative methods for linear systems
Seemingly minor variations in problem setting can lead to substantial variations in the complexity of the learning problem. Initially for the Jacobi method we posed a subspace regression problem as an approximation of leading eigenspaces of error propagation matrix $I-D^{-1}A$, where $D$ is diagonal of $A$. Neural networks with and without subspace embedding completely failed to learn. After inspection of the dataset we found that the leading eigenspace contains a complicated mixture of functions with low and high frequencies. Since the learning problem appeared to be completely hopeless, we reformulated subspace regression using error propagation matrix of damped Jacobi iteration $I-\omega D^{-1}A$ with $\omega=0.9$. In contrast to the standard Jacobi method, the damped version leads to subspaces formed by low frequency functions. As evident from Figure 3(b) the resulting mapping is easily learnable. A more detailed report can be found in Appendix H.
@@ Appendix A Proof of Theorem 1
    where $P_{A}=A\left(A^{\top}A\right)^{-1}A^{\top},P_{B}=B\left(B^{\top}B\right)^{-1}B^{\top}$ are orthogonal projectors on the columns spaces of $A$ and $B$. When QR decompositions $A=Q_{A}R_{A},B=Q_{B}R_{B}$ are available, projectors become $P_{A}=Q_{A}Q_{A}^{\top},P_{B}=Q_{B}Q_{B}^{\top}$ and identity (8) can be verified by algebraic manipulations
@@
    where $P_{A}=A\left(A^{\top}A\right)^{-1}A^{\top}$ is orthogonal projector on the columns space of $A$. Using $I=\left(I-P_{A}\right)+P_{A}$, and $A\left(I-P_{A}\right)=\left(I-P_{A}\right)A=0$ we obtain
@@
    The last equality holds since $P_{A}Q_{B}$ and $A$ share the same columns space. Given that $P_{A}$ does not depend on representative $A$ from $\lceil A\rceil$, and that $L_{2}(A,B;z)$ depends on $A$ only via $P_{A}$, we conclude that the same is true for $L_{2}(A,B;z)$.
@@ Appendix B Proof of Theorem 2
In most parts of the text we assumed working with the non-compact Stiefel manifold and in this theorem we have data on the compact Stiefel manifold (see [1] for definitions). We specify how one can compute $Q_{A}$ and $\dot{Q}_{A}$ having $A$ and $\dot{A}$. One may start from any stable version of Cholesky QR, e.g., [14], [37], and obtain
@@
where $y_{i}$ are columns of $Y$, $u_{i}$ are columns of $U$, $\sigma_{i}$ are diagonal elements of $\Sigma$, $\widetilde{\Sigma}=\Sigma-(\sigma_{1}-1)e_{1}e_{1}^{\top}$, that is, $\widetilde{\Sigma}$ can be obtained from $\Sigma$ by replacing $\Sigma_{11}=\sigma_{1}$ by $1$. Clearly $B(t)$ is geodesic and $\left\|\dot{B}(t)\right\|_{F}^{2}=\left\|\dot{A}(t)\right\|_{F}^{2}-\sigma_{1}<\left\|\dot{A}(t)\right\|_{F}^{2}$. Next we show that principal angles between $A(t)$ and $B(t)$ are all zero. To see this we observe that
@@ Appendix D Proof of Theorem 3
To illustrate this process, consider $E(z_{1},\dots,z_{D})=a_{1}z_{1}^{2}+a_{2}z_{2}^{2}$, where $a_{2}\gg a_{1}$. If we follow procedure outlined above we will see that first lattice points encountered are $(1,1),(2,1),(3,1),(4,1),\dots$. So for considered $a_{1},a_{2}$ the subspace of first $3$ eigenvectors is a span of $\phi_{1,1},\phi_{2,1},\phi_{3,1}$, and the eigenvector that appears on position $3$ is $\phi_{3}$. To describe the map from $a_{1},\dots,a_{D}$ to $\phi_{k}$ or $V_{k}$, this procedure needs to be repeated for all possible positive values of real coefficients $a_{1},\dots,a_{D}$.
From the algorithm above one can deduce that for given $a_{1},\dots,a_{D}$ the first time eigenvector with indices $i_{1},\dots,i_{D}$ appears in the sequence of eigenvectors is the first time ellipsoid crosses $i_{1},\dots,i_{D}$. The position $k$ of this eigenvector will be proportional to the normalised volume of the ellipsoid $V_{e}(a_{1},\dots,a_{D})\big/V_{s}$, where $V_{e}(a_{1},\dots,a_{D})$ is a volume of $D$ dimensional ellipsoid with semi-axes $a_{1},\dots,a_{D}$ and $V_{s}$ is a volume of $D$ dimensional sphere with radius $1$.
@@ Appendix E Monte Carlo experiments for eigenproblem with constant coefficients
To illustrate consequences of Theorem 3 we perform simple Monte Carlo experiment. For the case $D=3$, we generate $a_{1},a_{2},a_{3}$ from uniform distribution on $[0,1]$ repeatedly and record distinct subspaces of dimension $k=10$. We select unique colour for each subspace and draw them for each point $a_{1},a_{2},a_{3}$ on the plane perpendicular to $\begin{pmatrix}1&1&1\end{pmatrix}^{\top}$. This illustration appears in Figure 6(a). Theorem 3 suggest that the large number of distinct subspaces is a result of selection of $k$ vectors among a small number of candidates. This suggests we may decrease complexity of the function from coefficients to subspaces by predicting excessive number of vectors. In our experiments with neural networks this redundant mapping is learned, here we build the mapping with simple greedy strategy. In place of function $a_{1},a_{2},a_{3}\rightarrow V_{10}$ we consider $a_{1},a_{2},a_{3}\rightarrow V_{10}\cup\left\{v_{1},\dots,v_{k}\right\}$ where $v_{1},\dots,v_{k}$ are first $k$ most abundant eigenvectors. By appending additional vectors we decrease the number of distinct subspaces. The result of this greedy simplification appear in Figure 6(b) with $10$ additional vectors and in Figure 6(c) with $20$ additional vectors. As we see the number of distinct subspaces rapidly decreasing.
@@ F.1 Datasets
where $u,v$ are component of random normalised vector, $c=\sqrt{2}r_{e}$. Potentials $V_{1}$ and $V_{2}$ are i.i.d. with parameters: order of polynomial is $2$, $r_{e}$ is uniformly distributed on $[1,5]$, $d$ is uniformly distributed on $[10,40]$, coefficients of $\widetilde{q}$ are sampled from uniform distribution on $[0,3]$ and $c$ is sampled from uniform distribution on $[10,13]$. To discretise the problem we use finite difference and uniform $100\times 100$ grid on $7,7]^{2}$.
@@ F.2 Architectures and training
For $D=3$ grid search is not practical, so we select $N_{\text{layers}}=4$, $N_{\text{features}}=128$, $N_{\text{modes}}=16$, $\text{lr}=10^{-3}$, $\gamma_{\text{decay}}=0.5$, $N_{\text{decay}}=100$. The size of the train set is $800$, the size of test set is $200$. The number of epochs is $1000$. Architecture is training to approximate subspace spanned by first $3$ eigenvectors.
@@ G.2 Architectures and training details
DeepONet is used for standard regression, and the intrusive technique with bases extracted from branch net. As a branch net of DeepONet we select a classical convolution network with downsampling by factor of $2$ along each dimension and the increase of the number of hidden features by factor of $2$. Branch net is defined by the number of features after encoder $N_{e,b}$, kernel size of convolution $k_{b}$, and number of layers $N_{b}$. Trunk net is MLP which is defined by the number of hidden neurons $N_{f,t}$, number of layers $N_{t}$ and the size of basis on the output layer $N_{\phi}$. In out grid searches we used $N_{e,b}\in[4,5]$, $k_{b}\in[3,7]$, $N_{b}=4$, $N_{f,t}=N_{\phi}\in[100,200]$, $N_{t}\in[3,4]$.
@@ G.3 Additional results
We provide two additional results. For elliptic equations we compare optimality of learned bases. The results are in Figure 7(a) and Figure 7(b). Interestingly, global POD leads to a better basis than both FFNO and DeepONet. For DeepONet this is expected, since trunk net does not depend on parameters of PDE. For FFNO the result is more surprising, because the basis is extracted from the last hidden layer, so it explicitly depends on parameters. Bases of DeepPOD and subspace regression are the most optimal one, still they are highly non-optimal when compared with local POD. For the Burgers equation we compare DeepPOD and two variants of subspace regression in Table 6. SubReg(10)(10) is a subspace regression trained to approximate subspace spanned by first $10$ local POD basis functions and SubReg(5)(5) was trained with $5$ local POD basis functions. DeepPOD is an unsupervised method and was trained with the whole trajectory. This implies methods are sorted from left to right in the decrease of information they receive about solutions. Interestingly, the SubReg(5)(5) – method, learning the smallest subspace – performs better almost uniformly. A possible explanation is that an optimal subspace of dimension $5$ leads to good enough accuracy and is easier to learn than larger subspaces.

### v2

@@ 2.2 Examples of subspace regression problem
When equation (5) is discretised with finite difference or finite element method, it reduces to linear problem $A\phi=f$, where AA is large sparse matrix and $\phi,f$ are discretised solution and right-hand side of (5). To exploit sparsity of AA one can solve linear equation with relaxation method. General relaxation method split matrix additively $A=D+C$, where $D$ is regular with known inverse [27]. Given the split, if iteration scheme $x^{n+1}=x^{n}+D^{-1}\left(b-Ax^{n}\right)$ is convergent, steady state is exact solution $x=A^{-1}b$. Convergence is linear and its rate is defined by spectral radius of error propagation matrix $I-D^{-1}A$. To improve convergence rate, one can augment iterative method with coarse-grid correction [33]. This techniques allows one to remove influence of leading subspace $V$ of $I-D^{-1}A$ by solving small reduced linear system for error equation $V^{\top}AVe=r$, where $e$ and $r$ are error and residual in the subspace $\mathcal{S}(V)$. Naturally, subspace regression (1) for this problem approximates the mapping $A\rightarrow\mathcal{S}(V)$ or $k(x)\rightarrow\mathcal{S}(V)$ for linear systems resulting from equation (5).
Deflation for conjugate gradient. Krylov subspace methods provide a more systematic way to solve large sparse linear systems [27]. For linear system with symmetric positive definite matrix AA resulting from discretisation of equation (5), the method of choice is conjugate gradient (CG) [15]. Similar to other Krylov methods, on step $r$, CG identify optimal solution within Krylov subspace $\mathcal{K}_{r}=\text{span}\left\{b,Ab,\dots,A^{r-1}b\right\}$, where span refers to the subspace formed by linear combinations of vectors in the set. Since powers of AA are involved, the most readily available vectors are from the subspaces with large eigenvalues [28]. To improve convergence of method it is reasonably to include eigenspaces $V$ with small eigenvalues to the approximation space $\mathcal{K}_{r}$. The resulting method is deflated CG and the approximation space is $\mathcal{K}_{r}\cup\mathcal{S}(V)$ [29]. In analogy with previous example, the subspace regression problem considered approximates $A\rightarrow\mathcal{S}(V)$ or $k(x)\rightarrow\mathcal{S}(V)$, but this time $V$ spans eigenspaces with small eigenvalues of matrix AA.
@@ 3.1 Loss functions
    Loss function $L_{1}(A,B)=p-\left\|Q_{B}^{\top}Q_{A}\right\|_{F}^{2}$ satisfies requirements (2), where $A=Q_{A}R_{A},B=Q_{B}R_{B}$ are QR decompositions of AA and $B$, $\left\|\cdot\right\|_{F}$ is Frobenius norm.
@@
    Let $z\in\mathbb{R}^{k}$ be a random variable with zero mean and identity covariance matrix. Loss functions $L_{2}(A,B;z)=\min_{u}\left\|Au-Q_{B}z\right\|_{2}^{2}$ does not depend on the choice of AA from $\lceil A\rceil$, where $B=Q_{B}R_{B}$ is QR decomposition.
@@ 4.1 Eigenspace prediction
Loss without QRQR scales much better for larger subspace sizes. In Figure 1(a) we demonstrate wall clock training time for $L_{1}(A,B)$ and $L_{2}(A,B;z)$ (least squares problem is solved with normal equation) per epoch on the same hardware for the same FFNO architectures. For small subspace sizes the training time the methods are roughly on par, but with the increase of subspace size QR starts to drastically slow down training with $L_{1}$ loss.
@@
Loss $L_{2}(A,B;z)$ may become unstable. From results summarised in Figure 1(a) one can assume that $L_{2}(A,B;z)$ is always preferable. Results for $D=3$ problem (elliptic eigenproblem, grid $30\times 30\times 30$, prediction of first 33 eigenvectors) summarised in Table 2 reveal a more nuanced picture. Loss $L_{2}(A,B;z)$ clearly performs worse than $L_{1}(A,B)$ and even fails for subspace size $24$. The reason for that is numerical instability of solvers based on the normal equation. To stabilize $L_{2}(A,B;z)$ we apply Cholesky-QR2 [37]. The results for stabilised loss shows that accuracy becomes comparable to $L_{1}(A,B)$ and even slightly better for larger subspace dimensions.
@@ 4.3 Iterative methods for linear systems
Seemingly minor variations in problem setting can lead to substantial variations in the complexity of the learning problem. Initially for the Jacobi method we posed a subspace regression problem as an approximation of leading eigenspaces of error propagation matrix $I-D^{-1}A$, where $D$ is diagonal of AA. Neural networks with and without subspace embedding completely failed to learn. After inspection of the dataset we found that the leading eigenspace contains a complicated mixture of functions with low and high frequencies. Since the learning problem appeared to be completely hopeless, we reformulated subspace regression using error propagation matrix of damped Jacobi iteration $I-\omega D^{-1}A$ with $\omega=0.9$. In contrast to the standard Jacobi method, the damped version leads to subspaces formed by low frequency functions. As evident from Figure 3(b) the resulting mapping is easily learnable. A more detailed report can be found in Appendix H.
@@ Appendix A Proof of Theorem 1
    where $P_{A}=A\left(A^{\top}A\right)^{-1}A^{\top},P_{B}=B\left(B^{\top}B\right)^{-1}B^{\top}$ are orthogonal projectors on the columns spaces of AA and $B$. When QR decompositions $A=Q_{A}R_{A},B=Q_{B}R_{B}$ are available, projectors become $P_{A}=Q_{A}Q_{A}^{\top},P_{B}=Q_{B}Q_{B}^{\top}$ and identity (8) can be verified by algebraic manipulations
@@
    where $P_{A}=A\left(A^{\top}A\right)^{-1}A^{\top}$ is orthogonal projector on the columns space of AA. Using $I=\left(I-P_{A}\right)+P_{A}$, and $A\left(I-P_{A}\right)=\left(I-P_{A}\right)A=0$ we obtain
@@
    The last equality holds since $P_{A}Q_{B}$ and AA share the same columns space. Given that $P_{A}$ does not depend on representative AA from $\lceil A\rceil$, and that $L_{2}(A,B;z)$ depends on AA only via $P_{A}$, we conclude that the same is true for $L_{2}(A,B;z)$.
@@ Appendix B Proof of Theorem 2
In most parts of the text we assumed working with the non-compact Stiefel manifold and in this theorem we have data on the compact Stiefel manifold (see [1] for definitions). We specify how one can compute $Q_{A}$ and $\dot{Q}_{A}$ having AA and $\dot{A}$. One may start from any stable version of Cholesky QR, e.g., [14], [37], and obtain
@@
where $y_{i}$ are columns of $Y$, $u_{i}$ are columns of $U$, $\sigma_{i}$ are diagonal elements of $\Sigma$, $\widetilde{\Sigma}=\Sigma-(\sigma_{1}-1)e_{1}e_{1}^{\top}$, that is, $\widetilde{\Sigma}$ can be obtained from $\Sigma$ by replacing $\Sigma_{11}=\sigma_{1}$ by 11. Clearly $B(t)$ is geodesic and $\left\|\dot{B}(t)\right\|_{F}^{2}=\left\|\dot{A}(t)\right\|_{F}^{2}-\sigma_{1}<\left\|\dot{A}(t)\right\|_{F}^{2}$. Next we show that principal angles between $A(t)$ and $B(t)$ are all zero. To see this we observe that
@@ Appendix D Proof of Theorem 3
To illustrate this process, consider $E(z_{1},\dots,z_{D})=a_{1}z_{1}^{2}+a_{2}z_{2}^{2}$, where $a_{2}\gg a_{1}$. If we follow procedure outlined above we will see that first lattice points encountered are $(1,1),(2,1),(3,1),(4,1),\dots$. So for considered $a_{1},a_{2}$ the subspace of first 33 eigenvectors is a span of $\phi_{1,1},\phi_{2,1},\phi_{3,1}$, and the eigenvector that appears on position 33 is $\phi_{3}$. To describe the map from $a_{1},\dots,a_{D}$ to $\phi_{k}$ or $V_{k}$, this procedure needs to be repeated for all possible positive values of real coefficients $a_{1},\dots,a_{D}$.
From the algorithm above one can deduce that for given $a_{1},\dots,a_{D}$ the first time eigenvector with indices $i_{1},\dots,i_{D}$ appears in the sequence of eigenvectors is the first time ellipsoid crosses $i_{1},\dots,i_{D}$. The position $k$ of this eigenvector will be proportional to the normalised volume of the ellipsoid $V_{e}(a_{1},\dots,a_{D})\big/V_{s}$, where $V_{e}(a_{1},\dots,a_{D})$ is a volume of $D$ dimensional ellipsoid with semi-axes $a_{1},\dots,a_{D}$ and $V_{s}$ is a volume of $D$ dimensional sphere with radius 11.
@@ Appendix E Monte Carlo experiments for eigenproblem with constant coefficients
To illustrate consequences of Theorem 3 we perform simple Monte Carlo experiment. For the case $D=3$, we generate $a_{1},a_{2},a_{3}$ from uniform distribution on $[0,1]$ repeatedly and record distinct subspaces of dimension $k=10$. We select unique colour for each subspace and draw them for each point $a_{1},a_{2},a_{3}$ on the plane perpendicular to $\begin{pmatrix}1&1&1\end{pmatrix}^{\top}$. This illustration appears in Figure 6(a). Theorem 3 suggest that the large number of distinct subspaces is a result of selection of $k$ vectors among a small number of candidates. This suggests we may decrease complexity of the function from coefficients to subspaces by predicting excessive number of vectors. In our experiments with neural networks this redundant mapping is learned, here we build the mapping with simple greedy strategy. In place of function $a_{1},a_{2},a_{3}\rightarrow V_{10}$ we consider $a_{1},a_{2},a_{3}\rightarrow V_{10}\cup\left\{v_{1},\dots,v_{k}\right\}$ where $v_{1},\dots,v_{k}$ are first $k$ most abundant eigenvectors. By appending additional vectors we decrease the number of distinct subspaces. The result of this greedy simplification appear in Figure 6(b) with $10$ additional vectors and in Figure 6(c) with 2020 additional vectors. As we see the number of distinct subspaces rapidly decreasing.
@@ F.1 Datasets
where $u,v$ are component of random normalised vector, $c=\sqrt{2}r_{e}$. Potentials $V_{1}$ and $V_{2}$ are i.i.d. with parameters: order of polynomial is 22, $r_{e}$ is uniformly distributed on $[1,5]$, $d$ is uniformly distributed on $[10,40]$, coefficients of $\widetilde{q}$ are sampled from uniform distribution on $[0,3]$ and $c$ is sampled from uniform distribution on $[10,13]$. To discretise the problem we use finite difference and uniform $100\times 100$ grid on $7,7]^{2}$.
@@ F.2 Architectures and training
For $D=3$ grid search is not practical, so we select $N_{\text{layers}}=4$, $N_{\text{features}}=128$, $N_{\text{modes}}=16$, $\text{lr}=10^{-3}$, $\gamma_{\text{decay}}=0.5$, $N_{\text{decay}}=100$. The size of the train set is $800$, the size of test set is $200$. The number of epochs is $1000$. Architecture is training to approximate subspace spanned by first 33 eigenvectors.
@@ G.2 Architectures and training details
DeepONet is used for standard regression, and the intrusive technique with bases extracted from branch net. As a branch net of DeepONet we select a classical convolution network with downsampling by factor of 22 along each dimension and the increase of the number of hidden features by factor of 22. Branch net is defined by the number of features after encoder $N_{e,b}$, kernel size of convolution $k_{b}$, and number of layers $N_{b}$. Trunk net is MLP which is defined by the number of hidden neurons $N_{f,t}$, number of layers $N_{t}$ and the size of basis on the output layer $N_{\phi}$. In out grid searches we used $N_{e,b}\in[4,5]$, $k_{b}\in[3,7]$, $N_{b}=4$, $N_{f,t}=N_{\phi}\in[100,200]$, $N_{t}\in[3,4]$.
@@ G.3 Additional results
We provide two additional results. For elliptic equations we compare optimality of learned bases. The results are in Figure 7(a) and Figure 7(b). Interestingly, global POD leads to a better basis than both FFNO and DeepONet. For DeepONet this is expected, since trunk net does not depend on parameters of PDE. For FFNO the result is more surprising, because the basis is extracted from the last hidden layer, so it explicitly depends on parameters. Bases of DeepPOD and subspace regression are the most optimal one, still they are highly non-optimal when compared with local POD. For the Burgers equation we compare DeepPOD and two variants of subspace regression in Table 6. SubReg(10)(10) is a subspace regression trained to approximate subspace spanned by first $10$ local POD basis functions and SubReg(5)(5) was trained with 55 local POD basis functions. DeepPOD is an unsupervised method and was trained with the whole trajectory. This implies methods are sorted from left to right in the decrease of information they receive about solutions. Interestingly, the SubReg(5)(5) – method, learning the smallest subspace – performs better almost uniformly. A possible explanation is that an optimal subspace of dimension 55 leads to good enough accuracy and is easier to learn than larger subspaces.